# A Study of the Generalized Naming Game and Bayesian Naming Game as Dynamical Systems

### v1

@@ Abstract
We study the $\beta$-model ($\beta$-NG) and the Bayesian Naming Game (BNG) as dynamical systems. By applying linear stability analysis to the dynamical system associated with the $\beta$-model, we demonstrate the existence of a non-generic bifurcation with a bifurcation point $\beta_{c}=1/3$. As $\beta$ passes through $\beta_{c}$, the stability of isolated fixed points changes, giving rise to a one-dimensional manifold of fixed points. Notably, this attracting invariant manifold forms an arc of an ellipse. In the context of the BNG, we propose modeling the Bayesian learning probabilities $p_{A}$ and $p_{B}$ as logistic functions. This modeling approach allows us to establish the existence of fixed points without relying on the overly strong assumption that $p_{A}=p_{B}=p$, where $p$ is a constant.
@@ I Introduction
The naming game is a multi-agent model that aims to simulate the spontaneous emergence of consensus among agents who interact pairwise according to two possible processes: agreement and word/name learning Castellano et al. [2009]; Blythe [2015]; Baronchelli [2016]; Chen and Lou [2019]; Marchetti et al. [2020a]. In its simplest realization, the agents are allow to learn only two possible names, denoted here by the symbols A and B, according to some basic rules (we refer the reader to Refs. Baronchelli [2016]; Marchetti et al. [2020a] for their definitions). Furthermore, if the underlying topology on which the agents may lie, is neglected altogether Castello et al. [2009]; Marchetti et al. [2020a], we call such a model as to the minimal naming game (MNG).
The MNG model can be thought as a particular case of two variants: the generalized naming game ($\beta$-NG) Baronchelli et al. [2007] and the Bayesian naming game (BNG) Marchetti et al. [2020b, a, 2021]. Both variants add human features otherwise absent in agents of MNG. In $\beta$-NG, a parameter $\beta$ is introduced to take into account the probability of acknowledged influence Friedkin [1990], thus affecting the agreement processes, while in BNG the name learning processes are now dictated by the time-dependent probabilities $p_{A}\equiv p_{A}\left(t\right)$ and $p_{B}\equiv p_{B}\left(t\right)$ necessary for generalizing the concept $\mathcal{C}$ associated to words A and B, respectively  Tenenbaum [1999]; Tenenbaum and Griffiths [2001]; Murphy [2012]. It is worth noting here that such probabilities are a direct consequence of Bayes rule Jeffreys [1939]. Therefore, according to the latter model the agents become Bayesian-rational individuals that in an uncertain world find their best action using the laws of probability de Finetti [1989]; Sanborn and Chater [2016]. That being said, the minimal naming game can be recovered from the above two variants by setting their characteristic parameters, i.e. $\beta$ and $p_{A},p_{B}$, to unity as illustrated in cartoon of Fig. 1.
@@
In this paper, we shall pursue the latter approach, thereby finding novel interesting dynamical features that, to our best knowledge, have been overlooked so far.
[…]
In Section III, we shall address the issue related to the lack of spontaneous symmetry breaking for the dynamical system associated with BNG. This issue arises when attempting to find the solution by numerically integrating the respective first-order differential system, assuming the initial condition $I_{0}=\left(0.5,0.5\right)$ and setting $p_{A}=p_{B}=p$ where $p$ is a constant, simultaneously.
Denoting the solution as $\left(x,y\right)$, where $x\equiv x\left(t\right)$ and $y\equiv y\left(t\right)$ are the population fractions having either A or B in their inventories at time $t$ respectively, it is found that $\left(x,y\right)$ asymptotically converges to $\left(0.38,0.38\right)$. Thus, there is no convergence to consensus, similar to what occurs with MNG for the very same setup Castellano et al. [2009]; Marchetti et al. [2020a].
This is in contrast to what is found in multi-agent simulations, where consensus always emerges either at A or at B for MNG and BNG models. It has already been noted that such a discrepancy between simulations and dynamical systems arises from the fact that the dynamical system approach cannot account for the ubiquitous density fluctuations, or equivalently noise, present in numerical simulations Castellano et al. [2009]. However, this issue is particularly significant here, as the above setup was used to prove the existence of consensus for the dynamical system associated with BNG Marchetti et al. [2020b].
@@
To address this serious problem, we propose modeling the Bayesian probabilities as time-dependent logistic functions (see Eq. 12James et al. [2023]. Subsequently, by numerically integrating the modified equations, we demonstrate that consensus consistently emerges at either A or B. This outcome is attributed to the symmetry breaking induced by the time-dependent probabilities, disrupting the invariance under the swap of variables $x$ and $y$ (see Eqs. 10 and  11). Accordingly, the solution closely resembles that obtained from multi-agent simulations with system size $N\gtrsim 200$ Marchetti et al. [2020b].
Figure 1: The cartoon depicts the relationship between the MNG, $\beta$-NG, and BNG through their characteristic parameters: $\beta\in[0,1]$ and the learning probabilities $p_{A},p_{B}\in[0,1]$. Setting $\beta$ and $p_{A},p_{B}$ to unity in $\beta$-NG and BNG models, respectively, one recovers the MNG.
II The $\beta$-NG model as Dynamical System
@@
In mean-field approximation the time-evolution of $\beta$-NG model is dictated by the following system of first-order differential equations:
$\dot{x}=-xy+\beta\left(1-x-y\right)^{2}+Kx\left(1-x-y\right)\,,$ (1)
$\dot{y}=-xy+\beta\left(1-x-y\right)^{2}+Ky\left(1-x-y\right)\,,$ (2)
@@
where $\dot{x},\dot{y}$ the derivatives of $x,y$, with respect to time. We set $K\equiv\left(3\beta-1\right)/2$ in Eqs. 1, 2 while $\beta\in[0,1]$. Next, setting $\dot{x}=0$ and $\dot{y}=0$ into Eqs.1 and 2, it was previously determined that three isolated fixed points exist: A, B, and CBaronchelli et al. [2007]; Castellano et al. [2009]. Applying the LST revealed a first-order non-equilibrium transition at $\beta=\beta_{c}$. Furthermore, the system converges to point C for $\beta<\beta_{c}$, leaving a fraction of undecided agents equal to $1-2c_{\beta}$, while a spontaneous emergence of consensus occurs at either A or B for $\beta>\beta_{c}$. These findings are summarized in the first and third columns of Table 1.
@@
While the above analysis is certainly correct and has been confirmed by multi-agent simulations, it does not address the generic bifurcation that characterizes the dynamical system in question.
@@
In the following, we shall unravel the dynamical features of the dynamical system by applying the LSA to the problem at hand in a systematic way. First, let us define the functions $f_{1}\left(x,y\right),f_{2}\left(x,y\right)$ such that $\dot{x}=f_{1}\left(x,y\right),\dot{y}=f_{2}\left(x,y\right)$ according to Eqs. 1, 2. It is then possible to compute the eiggenvalues of the following Jacobian matrix $J$ evaluated at a given fixed equilibrium point with Cartesian coordinates $\left(\tilde{x},\tilde{y}\right)$ Strogatz [2018]
@@
In Eq. 4 we plot $\lambda_{3},\lambda_{4}$ as function of $\beta$ by means of Eq.. 4. From a direct comparison of the corresponding curves with those of $\lambda_{1},\lambda_{2}$, one immediately concludes that $C$ has the opposite equilibrium character of A and B for each each $\beta\in[0,1]\setminus\{\beta_{c}\}$, see Fig. 3. However, when $\beta=\beta_{c}$, $C$ belongs to the stable line as well because $\lambda_{3}=0$ while $\lambda_{4}$ negative (see Fig. 2).
@@
Overall, our analysis shows that the fixed points $A,B,C$ exist for all values of the parameter $\beta$ but there an exchange of stabilities between $A,B$ and $C$ as $\beta$ passes through $\beta_{c}$. In this regard, we can say that the dynamical system undergoes a generic bifurcation similar to that of the 1D differential equation discussed in Section I. For the sake of completeness, we summarize all these findings in Table 1.
@@
Next, the existence of a continuum of fixed points at the bifurcation point could be easily inferred by examining the streamline plot of the velocity field $\mathbf{v}=\left(\dot{x},\dot{y}\right)$ corresponding to Eqs. 1, 2. To this end, in Fig. 4 we plot such a velocity field on the model’s domain in Cartesian $xy$-plane: $\{(x,y)\in\mathbb{R}^{2}:0\leq x\leq 1,0\leq y\leq 1,0\leq x+y\leq 1\}$. It is then evident that there exists one-dimensional manifold of fixed points (orange solid line), and that all trajectories approach it along parallel lines. This finding is consistent with the stable line, that is, the negative negative half-axis of trace $\operatorname{tr}J$, obtained from LSA (see Fig. 2).
Figure 4: Streamline plot of $\mathbf{v}=\left(\dot{x},\dot{y}\right)$ corresponding to Eqs. 1, 2 at bifurcation point $\beta=\beta_{c}$. Here the stable line (orange solid line) corresponds to the graph of the map $x\mapsto 1/2\left(-\sqrt{3}\sqrt{4x-x^{2}}+x+2\right)$ with $x\in[0,1]$. The important initial condition $I_{0}=\left(0.5,0.5\right)$ is also shown.
@@
with determinant $\det\left(M\right)=3/4$. As a consequence its eigenvalues must the same sign, so Eq. 8 yield a conic section that is an ellipse with semi-major axis, semi-minor axis and eccentricity corresponding to $a=\sqrt{6}$, $b=\sqrt{2}$ and $e=\sqrt{2/3}$, respectively. The graph of such an ellipse with its center $0=\left(2,2\right)$ and foci $F_{1}=\left(2-\sqrt{2},2-\sqrt{2}\right)$ and $F_{2}=\left(2+\sqrt{2},2+\sqrt{2}\right)$ is shown in Fig. 5. Note that in such a case, point $C$ now has coordinates $\left(c_{\beta_{c}},c_{\beta_{c}}\right)$ with $c_{\beta_{c}}=2-\sqrt{3}$ and hence lies on the ellipse between points A and B.
@@
The paths with $\omega=1,2,3$ exhibit a pattern completely different from those at the bifurcation point, i.e., with $\omega=4,5,6$. While the former remain close to the sample mean $E[X_{t}]$ (black solid line) computed from $1,200$ trajectories, thus showing the typical mean-returning behavior of the Ornstein-Uhlenbeck (OU) process Uhlenbeck and Ornstein [1930]; Vasicek [1977]; Oksendal [2010], the latter seem to wander in a more random fashion. Preliminary statistical analysis Marchetti [prep] confirms that the sample paths at $\beta=0.2$ belong to OU process, while those at bifurcation point to a nonstandard Brownian motion Einstein [1905]; Langevin [1908]; Wiener [1923]; Ritort [2019].
Figure 6: Randomly chosen sample paths of $X_{t}$ ($X_{0}=0.5$) as functions of time $t$ from multi-agents simulations with $N=1,000$ agents. The paths with label $\omega=1,2,3$ and $\omega=4,5,6$ correspond to $\beta=0.2$ and $\beta=\beta_{c}$, respectively. Solid lines in black color and red color corresponds to sample mean $E[X_{t}]$ computed from $1,200$ trajectories for $\beta=0.2$ and $\beta=\beta_{c}$, respectively.
III The Bayesian Naming Game as a Dynamical System
According to the BNG model, agents must learn new, unknown names through their previous experiences and background knowledge. Furthermore, since the learning process occurs under uncertainty, a probabilistic approach is necessary for modeling such a process. In this context, a Bayesian framework proposed by Tenenbaum and co-workers seems appropriate Tenenbaum [1999]; Murphy [2012]; Griffiths and Tenenbaum [2006]; Griffiths and Kalish [2007]; Tenenbaum et al. [2011]; Perfors et al. [2011]; Lake et al. [2015]. As a result of this probabilistic framework, agents attempt to generalize by computing the learning probabilities $p_{A}$ and $p_{B}$, which depend on the positive examples stored by the agents during their pairwise interactions. In this regard, the BNG model paves the way for incorporating many human cognitive biases Tversky and Kahneman [1974]; Hahn [2014]; Sanborn and Chater [2016]; Ngampruetikorn and Stephens [2016]; Madsen et al. [2018] into the consensus dynamics, which were out of reach for the previous variants of naming game.

### v3

# Generalized Naming Game and Bayesian Naming Game as Dynamical Systems

@@ Abstract
We study the $\beta$-model ($\beta$-NG) and the Bayesian Naming Game (BNG) as dynamical systems. By applying linear stability analysis to the dynamical system associated with the $\beta$-model, we demonstrate the existence of a non-generic bifurcation with a bifurcation point $\beta_{c}=1/3$. As $\beta$ passes through $\beta_{c}$, the stability of isolated fixed points changes, giving rise to a one-dimensional manifold of fixed points. Notably, this attracting invariant manifold forms an arc of an ellipse. In the context of the BNG, we propose modeling the Bayesian learning probabilities $p_{A}$ and $p_{B}$ as logistic functions. This modelling approach allows us to establish the existence of fixed points without relying on the overly strong assumption that $p_{A}=p_{B}=p$, where $p$ is a constant.
@@ I Introduction
The naming game is a multi-agent model that aims to simulate the spontaneous emergence of consensus among agents who interact pairwise according to two possible processes: agreement and word/name learning Baronchelli et al. [2006]; Blythe [2015]; Baronchelli [2016]; Chen and Lou [2019]; Marchetti et al. [2020a]. In its simplest realization, the agents are allowed to learn only two possible names, denoted here by the symbols A and B. Furthermore, if the underlying topology on which the agents may lie, is neglected altogether Castello et al. [2009]; Marchetti et al. [2020a], we refer to such a model as the minimal naming game (MNG).
This multi-agent model can be thought as a particular case of two variants: the generalized naming game ($\beta$-NG) Baronchelli et al. [2007] and the Bayesian naming game (BNG) Marchetti et al. [2020b, a, 2021]. Both variants add human features otherwise absent in agents of MNG. In $\beta$-NG, a parameter $\beta\in[0,1]$ is introduced to take into account the probability of acknowledged influence Friedkin [1990], thus affecting the agreement processes, while in BNG the name learning processes are now dictated by the time-dependent probabilities $p_{A}\equiv p_{A}\left(t\right)$ and $p_{B}\equiv p_{B}\left(t\right)$ necessary for generalizing the concept $\mathcal{C}$ associated with words A and B, respectively  Tenenbaum [1999a]; Tenenbaum and Griffiths [2001]; Murphy [2012]. It is worth noting here that such probabilities are a direct consequence of Bayes rule Jeffreys [1939]. Therefore, according to the latter model the agents become Bayesian-rational individuals that in an uncertain world find their best action using the laws of probability de Finetti [1989]; Aumann [1976]; Aaronson [2005]; Sanborn and Chater [2016]. We refer the reader to Section II for a detailed description of MNG and its variants $\beta$-NG and BNG.
@@
In this paper, we shall pursue the latter approach, thereby finding novel interesting dynamical features of the dynamical systems associated with the $\beta$-NG and BNG, which, to our best knowledge, have been overlooked so far.
To begin with, by systematically applying the linear stability (LSA) to the dynamical system associated with $\beta$-NG we show that such a dynamical system undergoes to a non-generic bifurcation with bifurcation point $\beta=\beta_{c}$ where $\beta_{c}=1/3$ (see Section III). This is an important finding because it shows that the first-order (nonequilibrium) phase transition at $\beta=\beta_{c}$ obtained by means of numerical simulations in Ref. Baronchelli et al. [2007] corresponds to a special feature in the dynamical system approach: a bifurcation where several things happen simultaneously. In fact, there is an emergence of one-dimensional manifold of fixed points forming an arc of an ellipse, when $\beta$ goes through $\beta_{c}$. Additionally, this bifurcation affects the type of noise present in the trajectories from multi-agent simulations, as they are interpreted as sample paths of two important stochastic processes: the Brownian motion and Ornstein-Uhlenbeck (OU) process Erban et al. [2009]; Oksendal [2010].
@@
Regarding the dynamical system associated with the BNG, in Section IV, we shall show that the overly strong assumption that $p_{A}=p_{B}=p$, where $p$ is a constant, cannot be used to prove the existence of fixed points and, hence, the existence of consensus, as proposed in Ref. Marchetti et al. [2020b], in full generality, due to the lack of a spontaneous symmetry breaking process. To address this serious problem, we propose modeling the Bayesian probabilities as time-dependent logistic functions James et al. [2023]. Subsequently, by numerically integrating the corresponding differential system, it is shown that there exist two fixed points and also that its solution closely resembles that obtained from multi-agent simulations with a system size $N\gtrsim 200$ Marchetti et al. [2020b].
II Minimal Naming Game and its variants: $\beta$-NG and BNG
According to the BNG model, at each time-step, two agents are randomly chosen within a system of size $N$. Then, the first agent, who acts as a speaker, selects the name stored in her inventory or randomly chooses one of the two possible names A and B if both are therein stored, and then utters it to the second agent (the hearer)  [^1] In our simulations, the agents do not need to invent the names from scratch, as one name, either A or B, is assigned to each agent’s inventory at the beginning.. If the inventory of the hearer contains such a name, then both agents update their inventories, keeping only the conveyed name (agreement). Otherwise, the hearer learns the new name and subsequently adds it to her inventory (deterministic word learning, as it occurs with a probability equal to unity). Baronchelli et al. [2006]; Chen and Lou [2019]. Note that through the above pairwise interactions, the agents attempt to reach a consensus on the name to assign to a single object Baronchelli [2016]. Such a consensus can be achieved when all the agents have the same single name in their inventories.
According to the $\beta$-NG model, the previous agreement process requires a modification. Specifically, if both agents have the communicated name in their inventories, they are allowed to update their inventories with a probability of $\beta$, while nothing changes with a probability of $1-\beta$ Baronchelli et al. [2007].
On the other hand, the BNG model replaces the above deterministic learning process with a human-like word learning process under uncertainty based on Bayesian inference Tenenbaum [1999a, b]; Murphy [2012]; Griffiths and Tenenbaum [2006]; Griffiths and Kalish [2007]; Tenenbaum et al. [2011]; Perfors et al. [2011]; Lake et al. [2015], while leaving the agreement process unchanged. The key idea underpinning this model is that learning a word requires multiple cognitive efforts over an extended period of time, and also requires the learning (or generalization) of the concept $\mathcal{C}$ associated with it from a few positive examples (few-shot learning). In this regard, the BNG model paves the way for incorporating many human cognitive biases Tversky and Kahneman [1974]; Kahneman [2012]; Hahn [2014]; Sanborn and Chater [2016]; Ngampruetikorn and Stephens [2016]; Madsen et al. [2018] into the consensus dynamics, which were out of reach for the previous variants of naming game.
Accordingly, such a probabilistic model requires that the speaker not only utters a name but also, at the same time, provides a positive example associated with $\mathcal{C}$, randomly chosen from his or her inventory.  [^2] Therefore, the agents are now equipped with three inventories: one for storing the names A and/or B, and two for storing the positive examples associated with A and B.. For computational convenience, these positive examples are represented as points belonging to an axis-parallel rectangle, i.e., a rectangle concept, in $\mathbb{R}^{2}$  [^3] In our case the rectangle is defined the following tuple $\left(x,y,\sigma_{1},\sigma_{2}\right)$ where $x=0,y=0$ are the lower-left corner’s coordinates and $\sigma_{1}=3,\sigma_{2}=1$ the rectangle’s sizes along the $x$ and $y$-axis, respectively.. Next, the hearer needs to compute the Bayesian probability for generalizing $\mathcal{C}$ from the conveyed example together with those previously stored, assuming an Erlang prior Tenenbaum [1999a]; Marchetti et al. [2020b, 2021]. The learning process is successful whenever the value of such a probability is equal to or greater than a threshold value, which we set to $0.5$ Marchetti et al. [2020b].
In present work, all the multi-agent simulations will start with initial condition corresponding to $50\%$ of the agents knowing the word A and the remainder knowing the other word B Marchetti et al. [2020b]. Note that from a dynamical system point of view, this setup corresponds exactly to having the initial condition $I_{0}=\left(0.5,0.5\right)$ in the phase plane (see Fig. 3).
Furthermore, at the beginning of the simulation we randomly assigned an number of four positive examples to each agent’s inventory. However, there are different thresholds $n^{\ast}_{ex,A}=5$, and $n^{\ast}_{ex,B}=6$, of the number of positive examples stored by the agents, necessary for computing the probabilities for generalizing the concept, respectively. As a consequence of the last assumption, the words A and B are now distinguishable synonyms, and, at the same time, the learning of A is favored Marchetti et al. [2020b].
Finally, tt is worth noting here that in MNG and BNG models the consensus always emerges spontaneously, but the same can occurs only when $\beta>\beta_{c}$ in $\beta$-model. Furthermore, the MNG model can be recovered from its two variants by setting their characteristic parameters, i.e. $\beta$ and $p_{A},p_{B}$, to unity.
III The $\beta$-NG model as Dynamical System
@@
According to $\beta$-NG model, in mean-field approximation the time-evolution of the population fractions $x\equiv x\left(t\right)$ and $y\equiv y\left(t\right)$, having either A or B in their inventories at time $t$ respectively, is dictated by the following system of first-order differential equations Castello et al. [2009]:
$\dot{x}=-xy+\beta\left(1-x-y\right)^{2}+\frac{3\beta-1}{2}x\left(1-x-y\right)\,,$ (1)
$\dot{y}=-xy+\beta\left(1-x-y\right)^{2}+\frac{3\beta-1}{2}y\left(1-x-y\right)\,,$ (2)
where $\dot{x},\dot{y}$ the derivatives of $x,y$, with respect to time. Note that hereafter, for convenience, we will utilize the symbol $K\equiv\left(3\beta-1\right)/2$.
@@
Next, setting $\dot{x}=0$ and $\dot{y}=0$ into Eqs.1 and 2, it was previously determined that three isolated fixed points exist: $A=\left(1,0\right)$, $B=\left(0,1\right)$ and $C=\left(c_{\beta},c_{\beta}\right)$ whose coordinates in terms of parameter $\beta$ are given by $c_{\beta}=\left(5\beta+1-\sqrt{\Delta}\right)/4\beta$ where $\Delta\equiv 1+10\beta+9\beta^{2}$  Baronchelli et al. [2007]; Castellano et al. [2009]  [^4] Note that it was originally found that $\Delta=1+10\beta+17\beta^{2}$.. Additionally, the LST revealed a first-order non-equilibrium transition at $\beta=\beta_{c}$, but the corresponding bifurcation went unnoticed. In such a scenario, the system converges to point C for $\beta<\beta_{c}$, leaving a fraction of undecided agents equal to $1-2c_{\beta}$, while a spontaneous emergence of consensus occurs at either A or B for $\beta>\beta_{c}$. These findings are summarized in the first and third columns of Table 1.
@@
While the above analysis is certainly correct and has been confirmed by multi-agent simulations, it does not address the non-generic bifurcation that characterizes the dynamical system in question.
@@
In the following, we shall unravel the dynamical features of the dynamical system by applying the LSA to the problem at hand in a systematic way. First, let us define the functions $f_{1}\left(x,y\right),f_{2}\left(x,y\right)$ such that $\dot{x}=f_{1}\left(x,y\right),\dot{y}=f_{2}\left(x,y\right)$ according to Eqs. 1, 2. It is then possible to compute the eigenvalues of the following Jacobian matrix $J$ evaluated at a given fixed equilibrium point with Cartesian coordinates $\left(\tilde{x},\tilde{y}\right)$ Strogatz [2018]
@@
In Fig. 2 we plot $\lambda_{3},\lambda_{4}$ as function of $\beta$ by means of Eq.. 4. From a direct comparison of the corresponding curves with those of $\lambda_{1},\lambda_{2}$, one immediately concludes that $C$ has the opposite equilibrium character of A and B for each each $\beta\in[0,1]\setminus\{\beta_{c}\}$, see Fig. 2. However, when $\beta=\beta_{c}$, $C$ belongs to the stable line as well because $\lambda_{3}=0$ while $\lambda_{4}$ negative (see Fig. 1).
@@
[…]
@@
Next, the existence of a continuum of fixed points at the bifurcation point could be easily inferred by examining the streamline plot of the velocity field $\mathbf{v}=\left(\dot{x},\dot{y}\right)$ corresponding to Eqs. 1, 2. To this end, in Fig. 3 we plot such a velocity field on the model’s domain in Cartesian $xy$-plane: $\{(x,y)\in\mathbb{R}^{2}:0\leq x\leq 1,0\leq y\leq 1,0\leq x+y\leq 1\}$. It is then evident that there exists one-dimensional manifold of fixed points (orange solid line), and that all trajectories approach it along parallel lines. This finding is consistent with the stable line, that is, the negative negative half-axis of trace $\operatorname{tr}J$, obtained from LSA (see Fig. 1). Furthermore, we note that the flow of the above 1D differential equation is similar to the 2D flow along the attracting invariant manifold corresponding to the (orange) stable “line” in the bifurcation diagram (see Fig. 1) and to the (orange) curve in the 2D phase plane (see Fig. 3).
Figure 3: Streamline plot of $\mathbf{v}=\left(\dot{x},\dot{y}\right)$ corresponding to Eqs. 1, 2 at bifurcation point $\beta=\beta_{c}$. Here the curve (orange) of the attracting invariant manifold corresponds to the graph of the map $x\mapsto 1/2\left(-\sqrt{3}\sqrt{4x-x^{2}}+x+2\right)$ with $x\in[0,1]$. The important initial condition $I_{0}=\left(0.5,0.5\right)$ is also shown.
@@
with determinant $\det\left(M\right)=3/4$. As a consequence its eigenvalues must have the same sign, so Eq. 8 yield a conic section that is an ellipse with semi-major axis, semi-minor axis and eccentricity corresponding to $a=\sqrt{6}$, $b=\sqrt{2}$ and $e=\sqrt{2/3}$, respectively. The graph of such an ellipse with its center $0=\left(2,2\right)$ and foci $F_{1}=\left(2-\sqrt{2},2-\sqrt{2}\right)$ and $F_{2}=\left(2+\sqrt{2},2+\sqrt{2}\right)$ is shown in Fig. 4. Note that in such a case, point $C$ now has coordinates $\left(c_{\beta_{c}},c_{\beta_{c}}\right)$ with $c_{\beta_{c}}=2-\sqrt{3}$ and hence lies on the ellipse between points A and B.
@@
The paths with $\omega=1,2,3$ exhibit a pattern completely different from those at the bifurcation point, i.e., with $\omega=4,5,6$. While the former remain close to the sample mean $E[X_{t}]$ (black solid line) computed from $1,200$ trajectories, thus showing the typical mean-returning behavior of the OU process Uhlenbeck and Ornstein [1930]; Vasicek [1977], the latter seem to wander in a more random fashion. Preliminary statistical analysis Marchetti [prep] confirms that the sample paths at $\beta=0.2$ belong to OU process, while those at bifurcation point to a (nonstandard) Brownian motion Einstein [1905]; Langevin [1908]; Wiener [1923]; Ritort [2019] (see Section A for details).
Figure 5: Randomly chosen sample paths of $X_{t}$ ($X_{0}=0.5$) as functions of time $t$ from multi-agents simulations with $N=1,000$ agents. The paths with label $\omega=1,2,3$ and $\omega=4,5,6$ correspond to $\beta=0.2$ and $\beta=\beta_{c}$, respectively. Solid lines in black color and red color correspond to sample mean $E[X_{t}]$ computed from $1,200$ trajectories for $\beta=0.2$ and $\beta=\beta_{c}$, respectively.
IV The Bayesian Naming Game as a Dynamical System